Vibrational Energy Transfer of I_{2}(X, v=23 and v=43) in a Helium Free Jet Expansion

\underline{T. A. Van Marter}, W. G. Lawrence, and M. C. Heaven, Department of Chemistry, Emory University, Atlanta, GA 30322.

Collision induced vibrational energy transfer of highly excited I_{2}(X) was studied in a helium free jet expansion at about 5K. Stimulated emission pumping was used to prepare a specified vibrational state. Laser induced fluorescence of the D-X transition was then used to probe the initial and the collisionally populated states. Two vibrational levels (v=23, v=42) of I_{2}(X) were examined in this study. Preliminary analysis suggests a weak vibrational dependance of the relaxation rate constants as well as a near temperature independent cross section when compared to room temperature data. These findings are in disagreement with the predictions of the Landau-Teller model. A more detailed analysis will be performed and the results presented. Work supported by AFOSR under grant F49620-95-1-0010